A clinician wants to see how an anti-anxiety drug will effect panic disorder. In a population of patients diagnosed with a panic disorder, the average score is 36 on the State-Trait Anxiety Inventory (STAI). The clinician has 27 outside patients take the recommended dosage of the drug. After a month, the clinician tested the outside patients on STAI and obtains an average of 32 and a standard deviation of 14. What can the clinician conclude with an α of 0.05?

a) What is the appropriate test statistic?
---Select--- na z-test One-Sample t-test Independent-Samples t-test Related-Samples t-test

b)
Population:
---Select--- the drug outside patients anxiety panic disorder patients with panic disorder
Sample:
---Select--- the drug outside patients anxiety panic disorder patients with panic disorder

c) Compute the appropriate test statistic(s) to make a decision about H0.
(Hint: Make sure to write down the null and alternative hypotheses to help solve the problem.)
critical value =  ; test statistic =
Decision:  ---Select--- Reject H0 Fail to reject H0

d) If appropriate, compute the CI. If not appropriate, input "na" for both spaces below.
[  ,  ]

e) Compute the corresponding effect size(s) and indicate magnitude(s).
If not appropriate, input and/or select "na" below.
d =  ;   ---Select--- na trivial effect small effect medium effect large effect
r2 =  ;   ---Select--- na trivial effect small effect medium effect large effect

f) Make an interpretation based on the results.

Panic disorder of outside patients is significantly higher after taking the anti-anxiety drug.

Panic disorder of outside patients is significantly lower after taking the anti-anxiety drug.    

The anti-anxiety drug did not have a significant impact on panic disorder.
Answer #1

a) One-Sample t-test

b)

Population:  patients with panic disorder

Sample:outside patients

c)

Ho :   µ =   36                  
Ha :   µ ╪   36       (Two tail test)          
                          
Level of Significance ,    α =    0.05                  
sample std dev ,    s =    14.0000                  
Sample Size ,   n =    27                  
Sample Mean,    x̅ =   32.0000                  
                          
degree of freedom=   DF=n-1=   26                  
                          
Standard Error , SE = s/√n =   14.0000   / √    27   =   2.6943      
t-test statistic= (x̅ - µ )/SE = (   32.000   -   36   ) /    2.6943   =   -1.485
                          
critical t value, t* =    ±   2.0555   [Excel formula =t.inv(α/no. of tails,df) ]              
                          
Decision: test stat > -2.0555, Fail to reject null hypothesis                       

d)

Cohen's d=|(mean - µ )/std dev|=   0.29

r²=t²/(t²+df) = 0.078

f)

The anti-anxiety drug did not have a significant impact on panic disorder.